# Semantic Refinement with LLMs for Graph Representations

### v1

@@ 6 Conclusion
In this work, we proposed DAS, a data-centric framework for iterative, structure-aware node semantic refinement on graphs. By coupling a fixed GNN with a large language model through a model-conditioned history memory, DAS enables node semantics to be progressively adapted under joint structural context and task feedback. Experiments on both text-rich and text-free graphs show that DAS consistently improves over strong LLM-as-enhancer baselines in structure-dominated settings while remaining competitive in semantics-rich regimes. More broadly, this work highlights a new direction for graph learning, where input representations are treated as dynamic, task-adaptive states rather than static features.
Limitations
@@
Although DAS provides a flexible data-centric mechanism for adapting node semantics to varying graph regimes, several limitations remain. First, the approach incurs non-trivial computational overhead because each refinement round requires a separate LLM inference pass over all nodes. While our experiments suggest that a small number of refinement iterations suffice on medium-sized graphs, scaling DAS to very large graphs or to settings requiring repeated retraining would be challenging without further optimization or model compression strategies. To this end, we can modify its refinement process refining only uncertain or representative nodes, offering a potential path toward scalability on larger graphs. Second, the quality and faithfulness of refined descriptions depend on the underlying language model. Weaker LLMs may fail to preserve structural cues, and even strong LLMs can occasionally introduce factual inconsistencies. Although our prompts explicitly discourage hallucination and require strict adherence to the original information, semantic drift may still occur in some cases. Finally, our study focuses primarily on node classification. Whether the same refinement dynamics hold for tasks such as link prediction, clustering, or graph-level classification remains an open question and a promising area for future work.
Ethical Considerations
This work relies on large language models to generate or refine textual node descriptions. As with any LLM-based system, the outputs may inherit or amplify biases present in the pretraining data. When DAS is applied to graphs whose nodes represent people or sensitive entities, refined descriptions could potentially introduce or reinforce demographic stereotypes, even when the graph structure itself does not encode such information. Care should therefore be taken when deploying the method in sensitive domains. DAS also requires repeated LLM calls during iterative refinement. Depending on the model and API used, this may entail substantial computational and environmental cost. Although our experiments rely on a small number of refinement rounds, practitioners should be mindful of the carbon footprint and financial cost associated with large-scale deployment. Finally, DAS modifies node-level textual descriptions during training. While these refined descriptions are not intended for human consumption and are not used to produce new factual knowledge, they may nonetheless be misinterpreted as authoritative explanations of the underlying graph if presented without context. We therefore recommend that systems built on DAS avoid exposing refined descriptions directly to end users unless appropriate disclaimers are provided.
@@ A.1 GNNs with Fixed Inductive Bias
Classical graph representation learning is largely built upon fixed inductive biases that encode how structural and semantic information is propagated and aggregated over the graph. Early message-passing GNNs (kipf2017semisupervised; hamilton2017inductive; velickovic2018graph) propagate input node features through local neighborhood aggregation to learn task-specific representations. By design, these architectures favor locality and homophily (xu2018how; luan2022revisiting), thereby imposing a strong but fixed prior on how predictive signals are assumed to distribute over the graph. Complementary to message-passing models, structure-only methods characterize nodes by their positional or role similarity using random walks and structural homophily. Representative approaches include DeepWalk (perozzi2014deepwalk), node2vec (grover2016node2vec), and struc2vec (ribeiro2017struc2vec), which learn embeddings purely from graph topology without relying on node semantics. While highly effective in structure-dominated settings, these methods likewise rely on pre-specified structural assumptions that remain fixed across graph domains. Motivated by this limitation, a large body of subsequent work has sought to address structure–semantics mismatch primarily from the model side, by designing new architectures and learning mechanisms with enhanced inductive biases.
@@ A.2 Model-Centric Adaptation
A prominent line of model-centric approaches seeks to address structure–semantics heterogeneity by directly strengthening the inductive bias of GNN architectures (xue2024self; han2024sadenas). These methods go beyond standard message passing by encoding richer structural priors into the model design. For instance, maron2018invariant; chen2019equivalence; maron2019provably propose $k$-order Weisfeiler–Lehman (WL) GNNs to emulate the $k$-WL test within neural architectures, while murphy2019relational; loukas2019graph introduce positional and relational encodings to enhance representational power. Another line of work leverages random walk kernels to guide the message passing process, further enriching the inductive bias of existing GNNs (jin2022raw; tonshoff2021walking; wang2025beyond; wang2025generative). Despite improved expressivity, these models remain fundamentally model-centric: the inductive biases are still explicitly predefined by architecture design. Moreover, many remain theoretically bounded by the $k$-WL hierarchy (zhang2024beyond), suggesting that architectural enhancement alone cannot offer a principled solution to the open-ended diversity of real-world graphs.
@@
Another emerging model-centric paradigm treats LLMs as direct graph reasoners. These methods linearize graph structures and node attributes into natural language prompts and rely on the general reasoning capabilities of LLMs for training-free or lightly supervised graph classification and question answering (zhao2023graphtext; guo2023gpt4graph; wang2023can; kong2024gofa; chen2024llaga). For example, wang2023can describes graphs in natural language and applies LLMs to solve basic graph reasoning tasks, while GOFA (kong2024gofa) and LLaGA (chen2024llaga) operate over serialized graph representations or graph embeddings for downstream inference. By bypassing explicit message passing, these approaches effectively replace graph-specific inductive biases with the intrinsic reasoning priors of LLMs. However, this paradigm remains model-centric: structural information is processed solely according to the LLM and the serialization scheme, and context length limits together with the loss of explicit topology constrain scalability and long-range structural modeling.
@@
Beyond architectural modification and language-based reasoning, another class of model-centric approaches introduces auxiliary models to enrich the input representations of graph learners. Early work such as TextGCN (yao2019textgcn) demonstrates the benefit of incorporating external textual semantics into graph learning. More recently, with the advent of LLMs, a growing body of methods leverage LLMs as semantic enhancers to generate or refine node descriptions for downstream GNNs (chen2024text; zhang2024text; yan2023comprehensiveTAGreview; he2023harnessing; chen2024exploring; wang2024tans; yang2024latexgcl; fang2024gaugllm). In parallel, some works align auxiliary models with GNNs via joint training or embedding alignment (yang2021graphformers; zhao2022learningglem; wen2023augmenting2p2). Despite their effectiveness, these approaches remain model-centric: auxiliary models inject additional semantic or structural inductive biases, while the resulting node representations are typically treated as static inputs by the downstream GNN rather than being refined under task-driven feedback.
@@ A.3 Data-Centric Adaptation
A large body of work focuses on graph data augmentation, where node features or graph structures are perturbed to construct multiple views of the same graph for invariant representation learning (zhao2021data; zhu2021graph; suresh2021adversarial; you2020graph; you2021graph; wang2023heterogeneous; wang2024select). For example, GraphCL (you2020graph) applies a set of predefined structural and feature augmentations to generate contrastive graph views, enabling a model to capture augmentation-invariant information. These methods are effective for improving robustness and transferability, but the underlying node semantics are not explicitly refined toward task-specific semantic–structural alignment.
@@
Another line of data-centric work focuses on graph structure learning, which aims to optimize or reconstruct graph connectivity to better support GNN training (jin2020graph; liu2022towards; zhao2021heterogeneous; zhang2025mopi; perozzi2024let). These approaches adapt the graph topology by removing spurious edges or adding task-relevant connections, thereby modifying the structural substrate on which message passing operates. However, they primarily operate at the level of graph structure and do not directly model how node semantics should be adapted under different structure–semantics regimes.
@@
In addition, several studies explore pseudo-labeling and self-training schemes to guide representation learning in low-label settings (chen2023label). While effective for label efficiency, such methods treat node features as fixed inputs and do not address the problem of task-driven semantic adaptation under structural context.
The Position of Our Work.
In contrast to the above data-centric paradigms, our work targets a fundamentally different objective. Rather than augmenting data for invariance, modifying graph structure, or propagating pseudo-label supervision, we focus on task-driven refinement of node semantics themselves. Our method treats node semantics as adaptive variables that are progressively reshaped under structural context and predictive feedback from a downstream GNN, enabling direct handling of structure–semantics heterogeneity at the level where the balance between structure and semantics is instantiated.
Appendix B Understanding DAS: Why Iterative Semantic Refinement Works
We provide a theoretical interpretation of DAS as a generalized Majorization–Minimization (MM) procedure (sun2016majorization). We show that the iterative refinement loop monotonically decreases a task-adaptive surrogate objective, where the model-conditioned memory induces a tractable majorizer for semantic refinement.
B.1 Problem Formulation
Let $\mathcal{D}=\{d_{v}\}_{v\in\mathcal{V}}$ be the set of node descriptions and let $\mathbf{t}(d_{v})$ denote the sentence embedding of $d_{v}$. We define the global objective as
$\mathcal{J}(\theta,\mathcal{D})=\sum_{v\in\mathcal{V}_{\mathrm{train}}}\ell\!\left(y_{v},\,g_{\theta}(\mathbf{x}(d_{v}),G)\right)+\lambda\,\mathcal{R}(\mathcal{D}),$ (11)
where $\ell$ is the supervised loss and $\mathcal{R}$ is a memory-consistency regularizer defined below.
Implicit Memory-Consistency Regularizer.
At iteration $t$, DAS retrieves an exemplar set $\mathcal{S}_{v}^{(t)}$ from the model-conditioned memory $\mathcal{B}^{(t)}$. Given a candidate description set $\mathcal{D}$, consider the following admissible anchor set for each node:
$\mathcal{M}_{v}(\mathcal{D})$ $=\left\{\frac{1}{|\mathcal{S}_{v}|}\sum_{u\in\mathcal{S}_{v}}\mathbf{t}(d_{u})\right\},$ (12)
$\text{where }\;C$ $:\;\mathcal{S}_{v}\subseteq\mathcal{V}\setminus\{v\},\;|\mathcal{S}_{v}|\leq K.$
i.e., anchors induced by averaging embeddings of up to $K$ in-graph exemplars. In practice, $\mathcal{S}_{v}^{(t)}$ is selected by joint similarity and confidence filtering; the definition above abstracts this selection into a feasible set of anchors. We define
$\mathcal{R}(\mathcal{D})=\sum_{v\in\mathcal{V}}\min_{\mathbf{m}\in\mathcal{M}_{v}(\mathcal{D})}\left\|\mathbf{t}(d_{v})-\mathbf{m}\right\|_{2}^{2}.$ (13)
This regularizer encourages each node description to be close (in the embedding space) to a prototype summarizing reliable in-graph exemplars.
B.2 Memory-Induced Majorization
At iteration $t$, DAS fixes the retrieved exemplar sets $\{\mathcal{S}_{v}^{(t)}\}_{v\in\mathcal{V}}$ from $\mathcal{B}^{(t)}$ and defines the corresponding anchors
$\mathbf{m}_{v}^{(t)}=\frac{1}{|\mathcal{S}_{v}^{(t)}|}\sum_{u\in\mathcal{S}_{v}^{(t)}}\mathbf{t}\!\left(d_{u}^{(t)}\right).$ (14)
This yields the following surrogate regularizer:
$\Omega(\mathcal{D}\mid\mathcal{B}^{(t)})=\sum_{v\in\mathcal{V}}\left\|\mathbf{t}(d_{v})-\mathbf{m}_{v}^{(t)}\right\|_{2}^{2},$ (15)
and the iteration-$t$ surrogate objective
$\mathcal{U}(\mathcal{D}\mid\theta,\mathcal{B}^{(t)})=\sum_{v\in\mathcal{V}_{\mathrm{train}}}$ $\ell\!\left(y_{v},\,g_{\theta}(\mathbf{x}(d_{v}),G)\right)$
$+$ $\lambda\,\Omega(\mathcal{D}\mid\mathcal{B}^{(t)}).$ (16)
Lemma B.1 (Majorization and Tightness).
Fix $\theta$ and $\mathcal{B}^{(t)}$ constructed from $\mathcal{D}^{(t)}$. Then for any $\mathcal{D}$,
$\mathcal{J}(\theta,\mathcal{D})\leq\mathcal{U}(\mathcal{D}\mid\theta,\mathcal{B}^{(t)}),$ (17)
and the bound is tight at $\mathcal{D}=\mathcal{D}^{(t)}$, i.e., $\mathcal{J}(\theta,\mathcal{D}^{(t)})=\mathcal{U}(\mathcal{D}^{(t)}\mid\theta,\mathcal{B}^{(t)})$.
Proof.
By definition (13), for each $v$ and any feasible anchor $\mathbf{m}\in\mathcal{M}_{v}(\mathcal{D})$,
$\min_{\mathbf{m}^{\prime}\in\mathcal{M}_{v}(\mathcal{D})}\|\mathbf{t}(d_{v})-\mathbf{m}^{\prime}\|_{2}^{2}\leq\|\mathbf{t}(d_{v})-\mathbf{m}\|_{2}^{2}.$ (18)
In particular, using the feasible anchor $\mathbf{m}_{v}^{(t)}$ defined in (14) yields
$\min_{\mathbf{m}^{\prime}\in\mathcal{M}_{v}(\mathcal{D})}\|\mathbf{t}(d_{v})-\mathbf{m}^{\prime}\|_{2}^{2}\leq\|\mathbf{t}(d_{v})-\mathbf{m}_{v}^{(t)}\|_{2}^{2}.$ (19)
Summing over $v$ gives $\mathcal{R}(\mathcal{D})\leq\Omega(\mathcal{D}|\mathcal{B}^{(t)})$. Plugging into (11) proves (17). Tightness holds at $\mathcal{D}=\mathcal{D}^{(t)}$ because $\mathbf{m}_{v}^{(t)}$ is constructed from the current exemplar sets and current embeddings, hence it is one of the anchors considered by $\mathcal{R}$ at the current iterate. ∎
B.3 Monotonic Improvement of DAS
DAS alternates between (i) updating $\theta$ with $\mathcal{D}^{(t)}$ fixed, and (ii) refining $\mathcal{D}$ with $\theta$ and $\mathcal{B}^{(t)}$ fixed.
Theorem B.2 (Monotonic Descent).
Assume that at iteration $t$: (1) GNN step: the parameter update satisfies $\mathcal{J}(\theta^{(t+1)},\mathcal{D}^{(t)})\leq\mathcal{J}(\theta^{(t)},\mathcal{D}^{(t)}),$ e.g., by performing (approximate) descent on the supervised loss. (2) LLM refinement step: the refinement operator produces $\mathcal{D}^{(t+1)}$ such that $\Omega(\mathcal{D}^{(t+1)}\mid\mathcal{B}^{(t)})\leq\Omega(\mathcal{D}^{(t)}\mid\mathcal{B}^{(t)}).$ Then the global objective is monotonically non-increasing:
$\mathcal{J}(\theta^{(t+1)},\mathcal{D}^{(t+1)})\leq\mathcal{J}(\theta^{(t)},\mathcal{D}^{(t)}).$ (20)
Moreover, if $\mathcal{J}$ is bounded below, the sequence $\{\mathcal{J}(\theta^{(t)},\mathcal{D}^{(t)})\}$ converges to a finite limit.
Proof.
By Lemma B.1,
$\mathcal{J}(\theta^{(t+1)},\mathcal{D}^{(t+1)})\leq\mathcal{U}(\mathcal{D}^{(t+1)}\mid\theta^{(t+1)},\mathcal{B}^{(t)}).$ (21)
The refinement assumption implies
$\mathcal{U}(\mathcal{D}^{(t+1)}\mid\theta^{(t+1)},\mathcal{B}^{(t)})\leq\mathcal{U}(\mathcal{D}^{(t)}\mid\theta^{(t+1)},\mathcal{B}^{(t)}).$ (22)
By tightness in Lemma B.1, $\mathcal{U}(\mathcal{D}^{(t)}|\theta^{(t+1)},\mathcal{B}^{(t)})=\mathcal{J}(\theta^{(t+1)},\mathcal{D}^{(t)}).$ Combining the inequalities with the GNN step yields (20). The lower-boundedness follows from $\ell\geq 0$ and $\Omega\geq 0$. ∎
Theorem B.2 formalizes DAS as an MM procedure: $\mathcal{B}^{(t)}$ induces a majorizer by freezing the exemplar-induced anchors, and the LLM acts as a (possibly stochastic) descent oracle on the surrogate regularizer in the embedding space.

### v2

@@ 6 Conclusion and Limitations
In this work, we proposed GES, a data-centric framework for exemplar-guided, structure-aware node semantic refinement on graphs. By coupling a fixed GNN with a large language model through a model-conditioned memory, GES enables node semantics to be refined under joint structural context and task feedback. Experiments on both text-rich and text-free graphs show that GES consistently improves over strong LLM-as-enhancer baselines in structure-dominated settings while remaining competitive in semantics-rich regimes. More broadly, this work highlights a new direction for graph learning, where input representations are treated as dynamic, task-adaptive states rather than static features.
@@
Several limitations remain. First, GES requires a single LLM inference pass over all nodes, which incurs non-trivial overhead on large graphs; refining only uncertain or representative nodes offers a potential path toward scalability. Second, the quality of refined descriptions depends on the underlying language model—weaker LLMs may fail to preserve structural cues, and even strong LLMs can occasionally introduce factual inconsistencies despite explicit prompt constraints. Finally, our study focuses on node classification; whether the same refinement dynamics extend to link prediction, clustering, or graph-level tasks remains an open question for future work.
@@ A.1 GNNs with Fixed Inductive Bias
Classical graph representation learning is largely built upon fixed inductive biases that encode how structural and semantic information is propagated and aggregated over the graph. Early message-passing GNNs (Kipf and Welling, 2017; Hamilton et al., 2017; Veličković et al., 2018) propagate input node features through local neighborhood aggregation to learn task-specific representations. By design, these architectures favor locality and homophily (Xu et al., 2019; Luan et al., 2022), thereby imposing a strong but fixed prior on how predictive signals are assumed to distribute over the graph. Complementary to message-passing models, structure-only methods characterize nodes by their positional or role similarity using random walks and structural homophily. Representative approaches include DeepWalk (Perozzi et al., 2014), node2vec (Grover and Leskovec, 2016), and struc2vec (Ribeiro et al., 2017), which learn embeddings purely from graph topology without relying on node semantics. While highly effective in structure-dominated settings, these methods likewise rely on pre-specified structural assumptions that remain fixed across graph domains. Motivated by this limitation, a large body of subsequent work has sought to address structure–semantics mismatch primarily from the model side, by designing new architectures and learning mechanisms with enhanced inductive biases.
@@ A.2 Model-Centric Adaptation
A prominent line of model-centric approaches seeks to address structure–semantics heterogeneity by directly strengthening the inductive bias of GNN architectures (Xue et al., 2024; Han et al., 2024). These methods go beyond standard message passing by encoding richer structural priors into the model design. For instance, Maron et al. (2018); Chen et al. (2019); Maron et al. (2019) propose $k$-order Weisfeiler–Lehman (WL) GNNs to emulate the $k$-WL test within neural architectures, while Murphy et al. (2019); Loukas (2019) introduce positional and relational encodings to enhance representational power. Another line of work leverages random walk kernels to guide the message passing process, further enriching the inductive bias of existing GNNs (Jin et al., 2022; Tönshoff et al., 2021; Wang et al., 2025b; d). Despite improved expressivity, these models remain fundamentally model-centric: the inductive biases are still explicitly predefined by architecture design. Moreover, many remain theoretically bounded by the $k$-WL hierarchy (Zhang et al., 2024a), suggesting that architectural enhancement alone cannot offer a principled solution to the open-ended diversity of real-world graphs.
@@
Another emerging model-centric paradigm treats LLMs as direct graph reasoners. These methods linearize graph structures and node attributes into natural language prompts and rely on the general reasoning capabilities of LLMs for training-free or lightly supervised graph classification and question answering (Zhao et al., 2023a; Guo et al., 2023; Wang et al., 2023a; Kong et al., 2024; Chen et al., 2024a). For example, Wang et al. (2023a) describes graphs in natural language and applies LLMs to solve basic graph reasoning tasks, while GOFA (Kong et al., 2024) and LLaGA (Chen et al., 2024a) operate over serialized graph representations or graph embeddings for downstream inference. By bypassing explicit message passing, these approaches effectively replace graph-specific inductive biases with the intrinsic reasoning priors of LLMs. However, this paradigm remains model-centric: structural information is processed solely according to the LLM and the serialization scheme, and context length limits together with the loss of explicit topology constrain scalability and long-range structural modeling.
@@
Beyond architectural modification and language-based reasoning, another class of model-centric approaches introduces auxiliary models to enrich the input representations of graph learners. Early work such as TextGCN (Yao et al., 2019) demonstrates the benefit of incorporating external textual semantics into graph learning. More recently, with the advent of LLMs, a growing body of methods leverage LLMs as semantic enhancers to generate or refine node descriptions for downstream GNNs (Chen et al., 2024c; Zhang et al., 2024b; Yan et al., 2023; He et al., 2023; Chen et al., 2024b; Wang et al., 2025a; Yang et al., 2024; Fang et al., 2024). In parallel, some works align auxiliary models with GNNs via joint training or embedding alignment (Yang et al., 2021; Zhao et al., 2022; Wen and Fang, 2023). Despite their effectiveness, these approaches remain model-centric: auxiliary models inject additional semantic or structural inductive biases, while the resulting node representations are typically treated as static inputs by the downstream GNN rather than being refined under task-driven feedback.
@@ A.3 Data-Centric Adaptation
A large body of work focuses on graph data augmentation, where node features or graph structures are perturbed to construct multiple views of the same graph for invariant representation learning (Zhao et al., 2021b; Zhu et al., 2021; Suresh et al., 2021; You et al., 2020; 2021; Wang et al., 2023b; 2024a). For example, GraphCL (You et al., 2020) applies a set of predefined structural and feature augmentations to generate contrastive graph views, enabling a model to capture augmentation-invariant information. These methods are effective for improving robustness and transferability, but the underlying node semantics are not explicitly refined toward task-specific semantic–structural alignment.
@@
Another line of data-centric work focuses on graph structure learning, which aims to optimize or reconstruct graph connectivity to better support GNN training (Jin et al., 2020; Liu et al., 2022; Zhao et al., 2021a; Zhang et al., 2025; Perozzi et al., 2024). These approaches adapt the graph topology by removing spurious edges or adding task-relevant connections, thereby modifying the structural substrate on which message passing operates. However, they primarily operate at the level of graph structure and do not directly model how node semantics should be adapted under different structure–semantics regimes.
@@
In addition, several studies explore pseudo-labeling and self-training schemes to guide representation learning in low-label settings (Chen et al., 2023). While effective for label efficiency, such methods treat node features as fixed inputs and do not address the problem of task-driven semantic adaptation under structural context.
Appendix B Intuitive Motivation: Why Exemplar-Guided Refinement Produces Discriminative Embeddings
We provide a theoretical motivation for why exemplar-guided semantic refinement produces more discriminative node representations. The key insight is that the model-conditioned memory induces a task-adaptive prototype structure, and the LLM refinement operator moves node embeddings toward class-coherent prototypes defined by reliable exemplars.
B.1 Exemplar-Anchored Prototype Alignment
Let $\mathcal{D}=\{d_{v}\}_{v\in\mathcal{V}}$ be the set of node descriptions and let $\mathbf{t}(d_{v})$ denote the sentence embedding of $d_{v}$. For each node $v$, the model-conditioned memory $\mathcal{B}$ provides an exemplar set $\mathcal{S}_{v}$ of structurally aligned, semantically similar, and confidently classified nodes. Define the exemplar-induced prototype for node $v$ as
$\mathbf{m}_{v}=\frac{1}{|\mathcal{S}_{v}|}\sum_{u\in\mathcal{S}_{v}}\mathbf{t}(d_{u}).$ (3)
Because the exemplar set $\mathcal{S}_{v}$ is filtered by model confidence (low predictive entropy), nodes within $\mathcal{S}_{v}$ tend to be reliably classified. When these exemplars share the same true class as vv—which is encouraged by joint semantic–structural similarity—the prototype $\mathbf{m}_{v}$ approximates a class-coherent centroid in the embedding space.
The LLM refinement operator $d_{v}^{\prime}=\mathcal{M}(d_{v},\mathcal{S}_{v})$ reconstructs the node description by emphasizing discriminative cues from the exemplar set. In the embedding space, this operation can be interpreted as moving $\mathbf{t}(d_{v}^{\prime})$ closer to the prototype $\mathbf{m}_{v}$, effectively reducing the within-class embedding variance while preserving between-class separation. Formally, if the refinement satisfies
$\|\mathbf{t}(d_{v}^{\prime})-\mathbf{m}_{v}\|_{2}\leq\|\mathbf{t}(d_{v})-\mathbf{m}_{v}\|_{2},$ (4)
then the global memory-consistency objective
$\mathcal{R}(\mathcal{D})=\sum_{v\in\mathcal{V}}\|\mathbf{t}(d_{v})-\mathbf{m}_{v}\|_{2}^{2}$ (5)
is non-increasing after refinement.
B.2 Connection to Embedding Discriminability
The prototype alignment view predicts that GES refinement should increase the discriminability gap—defined as the difference between average same-class and different-class cosine similarities in the embedding space. We verify this prediction empirically in Section F.1: GES descriptions exhibit a 17–36% higher discriminability gap than TANS on citation graphs and a 26.4% higher inter/intra cluster ratio on airport graphs. Moreover, the discriminability gap is strongly correlated with downstream accuracy (Pearson $r=0.89$, $p<0.0001$), confirming that exemplar-guided refinement systematically improves the class-separability of node embeddings.